OpenStudy (anonymous):

whats the y intercept and x intercept of y=1/4x-1/3 show all your work

OpenStudy (anonymous):

I assume you mean y = (1/4)x -(1/3) Then y-intercept is where x = 0, so y-int = -(1/3) x-intercept is where y = 0, so x-int given by (1/4)x = (1/3), or x = 4/3